OpenStudy (anonymous):

reduce to lowest terms 210/294

OpenStudy (anonymous):

just take it easy, step by step... 105/147 = 15/21 = 5/7

OpenStudy (anonymous):

THANKS! I SEE WHERE I WENT WRONG!

OpenStudy (anonymous):

YOU ARE WELCOME! THANK YOU FOR THE MEDAL!!!